OpenStudy (anonymous):

Which equation best describes Tom’s age in the following: John is five years older than Tom. John is eighteen years old. 5 T equals eighteen t plus 5 equals eighteen eighteen minus t equals 5 T divided by 5 equals eighteen

OpenStudy (anonymous):

t + 5 = 18

OpenStudy (texaschic101):

j = t + 5 j = 18 18 = t + 5 18 - 5 = t 13 = t I agree...18 = t + 5 (and tom is 13)